# On-Device Emoji Classifier Trained with GPT-based Data Augmentation for a Mobile Keyboard

### v1

@@ 1 Introduction
  • We deployed the proposed emoji classifier as part of SwiftKey. (Section 4.3, 4.4)
The paper unfolds as follows: Section 2 introduces related work to emoji classification. Section 3 describes the proposed on-device emoji classifier with the data augmentation and reranking methods. This is followed by experimental results for this classifier on 3 different datasets and live experiments in Section 4. Finally, Section 5 concludes the paper.
@@ 2 Related Work
Different relevant work streams in the literature mitigated both on-device constrains and emoji classification challenges, which could be divided into two classes: multi-label emoji classification and single-label emoji classification.
@@
The first class focuses on predicting combination of emojis based on the input text Lin et al. (2019); Peng and Zhao (2021); Lee et al. (2022). For example, Peng and Zhao (2021) proposed an encoder-decoder model to predict combinations of 32 emojis. Lee et al. (2022) suggested a different approach by introducing an LSTM architecture to predict combinations of 64 emojis and explored the correlation between emoji and emotion, which showed better performance than Peng and Zhao (2021). However, they did not consider numerous emojis in their classifications. Peng et al. (2023) showed a successful approach to translate an input text to a combination of emojis while using synthetically generated samples from GPT. However, they did not consider the on-device low-latency and memory constrains for the modelling part. In addition, Kumar et al. (2021) presented another method to detect emoji boundary in input text and proposed an LSTM based architecture to predict combination of emojis. In Section 3.1, we will show that the Bert based architectures perform better than LSTM.
@@
The goal of the second class is single-emoji classification from the input text. For instance, Barbieri et al. (2017) proposed a Bidirectional Long Short-term Memory Networks (BLSTMs) model to classify among 20 most frequent emojis. Along the same line, Barbieri et al. (2018c); Felbo et al. (2017) introduced an LSTM architecture with label-wise attention targeting the SemEVal 2018 Shared Task Barbieri et al. (2018b) and another dataset with at most 200 emojis. The paper for SemEVal task is for predicting the most likely emoji out of 20 emojis to be used along with a tweet, which did not report results for the transformer architectures. Barbieri et al. (2020); Gamal et al. (2023) introduced emoji classifiers via transformers, but did not consider on-device latency and memory constrains. These constrains were not also covered by (Gandhi et al., 2022) in the Bert-based emoji classifier trained to predict 700 emojis with traditional sampling and cost-sensitive training techniques to account for imabalanced distribution of emoji dataset. Furthermore, Choudhary et al. (2018) proposed Siamese networks and focused on the performance of emoji classification of low-resource languages. Ramaswamy et al. (2019) successfully implemented an on-device emoji classifier for GBoard’s emoji prediction using an LSTM architecture, which is relatively low in computational requirements. Results in this paper were for English with only 100 emoji classes, which are not numerous.
Building on top of the literature, we focus on single-emoji classification and propose an on-device emoji classifier for SwiftKey that works relatively well on the imabalanced emoji dataset distribution especially when the number of classes are numerous, user favorites need to be considered, and latency/memory requirements are necessary.
@@ 3.1 Model Selection and Quantization
This observation is also confirmed when we train both an LSTM and Bert architectures on the 20-class SemEval dataset for 30 epochs with the AdamW optimizer and 1e-4 learning rate Gamal et al. (2023). SemEval dataset is a widely-used dataset for emoji classification benchmarking and competition that consists of 1.3 million examples of sentences and their corresponding emoji ground-truth. Our experiments along with results shared in Barbieri et al. (2018a) show that the top-1 F1 score and top-3 F1 score for the Bert and LSTM models on SemEval are 38%, 30% and 60%, 52%, respectively, which confirms that Bert-like and particulary MobileBert is eligibilble for the on-device emoji classifier (See Table 3 in Barbieri et al. (2018b)).
@@
To further decrease the size for the MobileBert on-device classifier while assessing its performance on larger classes, we train MobileNet for 5 epochs while taking one layer, 2 layers, and 3 layers from the architecture on another 100-class mixture of tweets (see Table 2). MobileBert with 2 layers shows a good accuracy performance at a relatively small runtime memory of 11MB. According to our simulations, this accuracy performance is insignificantly impacted by the int8 quantization while reducing the size further from 11MB to 3MB. With these results, we declare MobileBert with 2 layers as our proposed on-device emoji classifier that we serve in ONNX format for compatibility and ease of deployment developers (2021).
@@ 3.4 Emoji Classifier Reranking by User Favorites
The probability $P_{favorites}(emoji)$ is calculated as the number of times an emoji has been inserted divided by the total number of insertion events. $N_{favorites}$ denotes the number of unique emojis that have been inserted. The parameter $s$, empirically set to 4, is used to lessen the influence of favorites when the number of unique emoji insertions is low.
Equation 1 enables the favorites probability distribution to re-rank the output of the emoji model, thus adapting to each user’s preferences. As $\alpha$ increases, the impact of the favorites probability decreases while the influence of the emoji model increases, and vice versa.
@@ 4.1 Offline Accuracy Comparisons
Tables 3 and 4 show the performance for all models under test. We can observe that increasing the number of classes for the on-device emoji classifier lead to increasing the evaluation metrics because there is more coverage for the emoji classes. For example, the 1090C model has a Hit@24 of 10.7% better than the 90C model. These results go together with the increased coverage shown in Figure 2.
@@ 4.3 Live Experiments
As shown in Table 8, we observed that our 1090C model with augmentation lead to increased CTR, slightly less external emoji rate, and improved precision of 12%, 0.7%, and 4%, respectively relative to the logistic regression baseline. This performance is owing to the effectiveness of the 1090C model with augmentation and favorites reranking.
@@ 5 Conclusion
This paper presents an on-device emoji classifier deployed for SwiftKey trained with a new data augmentation method based on GPT that accounts for imbalanced emoji dataset distribution in numerous emoji classes. This proposed on-device emoji classifier with GPT’s data augmentation and method to adapt classifier output based on user preference increases the user engagement on SwiftKey’s predictive panel and performs well particularly on rare emojis while keeping the model size and inference time intact to run on-device. These results show ways to alleviate the emoji imbalanced dataset distribution. In the future, we wish to extend this classifier to predict multiple emojis.
Limitations
Our experiments show positive performance improvements for only English US, but ignores other languages. We wish to explore in our subsequent work how the on-device classifier and methods proposed in this paper scale with high and low resource languages. Low-resource languages, in particular, present a fascinating area of exploration due to the limited availability of samples.

### v2

@@ 1 Introduction
  • Deploy the emoji classifier in SwiftKey English with $\approx$10CTR increase. (Section 4.3)
Section 2 is the related work to emoji classification. Section 3 describes the on-device emoji classifier with the data augmentation and reranking methods. Experimental results for this classifier on 3 different datasets and live experiments are in Section 4. Section 5 concludes the paper.
@@ 2 Related Work
Different relevant work streams in the literature mitigated both on-device constrains and emoji classification challenges, which could be divided into two classes: output a combination of emojis at once and output a ranked list of emojis based on probability.
@@
The first class focuses on predicting combination of emojis based on the input text Lin et al. (2019); Peng and Zhao (2021); Lee et al. (2022). For example, Peng and Zhao (2021) proposed an encoder-decoder model to predict combinations of 32 emojis all at once. Lee et al. (2022) suggested a different approach by introducing an LSTM architecture to predict combinations of 64 emojis and explored the correlation between emoji and emotion, which showed better performance than Peng and Zhao (2021). However, they did not consider numerous emojis in their classifications. Peng et al. (2023) showed a successful approach to translate an input text to a combination of emojis while using synthetically generated samples from GPT. However, they did not consider the on-device low-latency and memory constrains for the modelling part. In addition, Kumar et al. (2021) presented another method to detect emoji boundary in input text and proposed an LSTM based architecture to predict combination of emojis. In Section 3.1, we will show that the Bert based architectures perform better than LSTM.
@@
The goal of the second class is to output a ranked list of emojis based on probability from the input text. For instance, Barbieri et al. (2017) proposed a Bidirectional Long Short-term Memory Networks (BLSTMs) model to classify among 20 most frequent emojis. Along the same line, Barbieri et al. (2018c); Felbo et al. (2017) introduced an LSTM architecture with label-wise attention targeting the SemEVal 2018 Shared Task Barbieri et al. (2018b) and another dataset with at most 200 emojis. The paper for SemEVal task is for predicting the most likely emoji out of 20 emojis to be used along with a tweet, which did not report results for the transformer architectures. Barbieri et al. (2020); Gamal et al. (2023) introduced emoji classifiers via transformers, but did not consider on-device latency and memory constrains. These constrains were not also covered by (Gandhi et al., 2022) in the Bert-based emoji classifier trained to predict 700 emojis with traditional sampling and cost-sensitive training techniques to account for imabalanced distribution of emoji dataset. Furthermore, Choudhary et al. (2018) proposed Siamese networks and focused on the performance of emoji classification of low-resource languages. Ramaswamy et al. (2019) successfully implemented an on-device emoji classifier for GBoard’s emoji prediction using an LSTM architecture, which is relatively low in computational requirements. Results in this paper were for English with only 100 emoji classes, which are not numerous for industry setting.
Building on top of the literature, this paper outputs a ranked list of likely emojis and focuses the novelty on: 1) On-device emoji classifier deployed for SwiftKey with large user base, latency/memory being considered; 2) Procedure for emojis class imbalance; 3) Relatively strong performance/analysis on imbalanced and numerous set of emojis tested in an industry setting; 4) Incorporation of dynamic user preferences for personalization.
@@ 3.1 Model Selection and Quantization
This observation is also confirmed when we train both an LSTM and Bert architectures on the 20-class SemEval dataset for 30 epochs with the AdamW optimizer and 1e-4 learning rate Gamal et al. (2023). SemEval dataset is a widely-used dataset for emoji classification benchmarking and competition that consists of 1.3 million examples of sentences and their corresponding emoji ground-truth. Our experiments along with results shared in Barbieri et al. (2018a) show that the top-1 F1 score and top-3 F1 score for the Bert and LSTM models on SemEval are 38%, 30% and 60%, 52%, respectively. We tested GPT3.5 on SemEval in zero-shot and saw 16% top-1 F1 score. This motivates that Bert-like and particularly MobileBert eligibility for the on-device emoji classifier (See Table 3 in Barbieri et al. (2018b)).
@@
To further decrease the size for the MobileBert on-device classifier while assessing its performance on larger classes, we train MobileNet for 5 epochs while taking one layer, 2 layers, and 3 layers from the architecture on another 100-class mixture of tweets (see Table 2). MobileBert with 2 layers shows a good accuracy performance at a relatively small runtime memory of 11MB. According to our simulations, this accuracy performance is insignificantly impacted by the int8 quantization while reducing the size further from 11MB to 3MB. With these results, we declare MobileBert with 2 layers as our proposed on-device emoji classifier that we serve in ONNX format for compatibility and ease of deployment developers (2021). To put runtime memory under control, an ONNX library was compiled only given the emoji classifier operators.
@@ 3.4 Emoji Classifier Reranking by User Favorites
The probability $P_{favorites}(emoji)$ is calculated as the number of times an emoji has been inserted divided by the total number of insertion events. $N_{favorites}$ denotes the number of unique emojis that have been inserted. $s$, empirically set to 4, is used to lessen the influence of favorites when the number of unique emoji insertions is low.
Equation 1 enables the favorites probability distribution to re-rank the output of the emoji model, thus adapting to each user’s preferences. As $\alpha$ increases, the impact of the favorites probability decreases while the influence of the emoji model increases, and vice versa. For rare emojis that are not part of the classifier’s output, $P_{model}(emoji|m)$ is set to 1e-4, ensuring that user’s frequently typed emojis still over time appear in predictions and alleviates the model retraining for new classes.
@@ 4.1 Offline Accuracy Comparisons
Tables 3 and 4 show the performance for all models under test. We can observe that increasing the number of classes for the on-device emoji classifier lead to increasing the evaluation metrics because there is more coverage for the emoji classes. For example, the 1090C model has a Hit@24 of 10.7% better than the 90C model. These results go together with the increased coverage shown in Figure 2. We also use D1 and phi3 zero-shot and saw that phi3 outputs meaningful emojis but may not be necessarily the user’s emoji choice leading to hit1 rates of 8% plus our models are much smaller.
@@ 4.3 Live Experiments
As shown in Table 8, our 1090C model with augmentation leads to increased CTR, less external emoji rate, and improved precision of 12%, 0.7%, and 4%, respectively relative to the logistic regression baseline. This performance is owing to the effectiveness of the 1090C model with augmentation and favorites reranking. Disregarding user preference personalization results in a 7% drop in CTR recall, highlighting its importance.
@@ 5 Conclusion
This paper presents an on-device emoji classifier for SwiftKey trained with a new data augmentation method based on GPT that accounts for imbalanced emoji dataset distribution in numerous emoji classes. This classifier with GPT’s data augmentation and method to adapt classifier output based on user preference increases the user engagement on SwiftKey’s predictive panel and performs well particularly on rare emojis while keeping the model size and inference time intact to run on-device. These results show ways to alleviate the emoji imbalanced dataset distribution. Although emojis are commonly used in English according to internal benchmarks, we aim to extend this classifier to other languages and rich media.